Please help me solve this question.
Find the equation of the hyperbola centered at (2,-3), whose slopes of asymptotes are y=± 3 and which opens up and down.

Answer by stanbon(75887)

Find the equation of the hyperbola centered at (2,-3), whose slopes of asymptotes are y=± 3 and which opens up and down.
-------------
Form:
-[(x-h)^2/a^2]+[(y-k)^2/b^2] = 1
-----------------------------------------
-[(x-2)^2/b^2]+[(y+3)^2/a^2 = 1
----
Asymptotes: y = 3x and y = -3x
Find a and b
a/b = 3
a = 3b
----
Note: you need to know a or b to finish
the hyperbola equation form.
You can get a or b if you know one of the vertices.
--------------------------------
Cheers,
Stan H.